# Learning to Reason for Long-Form Story Generation

### v1

@@ 6 Reinforcement Learning for Next-Chapter Prediction
We use the thresholded VR-CLI reward paradigm introduced in Equation (5) to define our reward for each (story-information, chapter, reasoning) tuple $(SI_{i},c_{i+1},\hat{p})$. Note that our ‘answer’ is the detailed plan $\hat{p}$ our policy model produces after reasoning over the story information. Example reasoning traces are in Appendix G. We use the reference model for our policy as our story-generator,  $\pi^{\mathcal{G}}$. We define our thresholded reward as:
$R(SI_{i},c_{i+1},\hat{p})=\begin{cases}0,&I(SI_{i},c_{i+1},\hat{p})<0.05\\ 0.5,&0.05\leq I(SI_{i},c_{i+1},\hat{p})<1\\ 0.9,&\text{if }1\leq I(SI_{i},c_{i+1},\hat{p})<2\\ 1,&\text{if }I(SI_{i},c_{i+1},\hat{p})\geq 2\\ \end{cases}$ (7)
@@
Our task and reward setup are relatively RL-algorithm agnostic; one could use an offline reasoning-reward dataset or any online policy learning algorithm that works with a verifiable reward. We hypothesize that online methods are preferable for NCP as reasoning styles change over training and preliminary tests show low reward prior to training: initially average improvement is -0.06% with Qwen 2.5 7B.
@@ 7.1 Model Comparisons
Our proposed method, training a policy model with VR-CLI and GRPO. Our story-generator is not further trained, but our policy model is trained to produce useful plans via our VR-CLI reward (Section 6): $\hat{p}\leftarrow\pi^{\mathcal{R}}_{\theta}(SI_{i});\ \ \hat{c}_{i}\leftarrow\pi^{\mathcal{G}}(SI_{i},\hat{p})$.
@@
We also compare to a model which attempts to predict the next chapter by first generating a reasoning trace; we do not train our story-generator or our reasoning model (they are both set to the same): $\hat{p}\leftarrow\pi^{\mathcal{R}}(SI_{i});\ \ \hat{c}_{i}\leftarrow\pi^{\mathcal{G}}(SI_{i},\hat{p})$.
@@
We finetune our story-generator on the NCP task with the SFT objective on next chapters. The generator predicts chapters like Base: $\hat{c}_{i}\leftarrow\pi^{\mathcal{G}}_{\theta}(SI_{i})$.
@@ 7.2 Human Evaluation
We evaluate the generated story continuations by eliciting pairwise preferences following evidence suggesting that relative judgments can be more reliable than absolute ones (Louviere et al., 2015; Stewart et al., 2005; Liu et al., 2024). We draw on the criteria outlined in previous work (Huot et al., 2024; Chakrabarty et al., 2024; Chhun et al., 2022) to evaluate the continuations along the following dimensions: (1) Plot: Does it exhibit events and turns that move the plot forward logically? (2) Creativity: Does the continuation have engaging characters, themes, and imagery, and avoid overly cliched characters and storylines? (3) Development:   Does it introduce characters and settings with appropriate levels of detail and complexity? (4) Language Use:   Is the language varied and rich, exhibiting rhetorical, linguistic, and literary devices? (5) Characters: Does it feature believable and conceptually consistent characters, including reasonable character arcs and development? (6) Overall Preference: Which of the two continuations did you prefer? Full instructions and recruitment details are in Appendix C. Annotators were recruited through Prolific if their primary language was English and their employment role was in creative writing.
To validate our annotation procedure we compute Fleiss’ kappa across a dataset of Base vs. Base-Reasoning comparisons. We find fair agreement across annotation dimensions, with highest agreement in Creativity and Overall Quality. More details are in Appendix D. Initial experiments showed gold-continuations were almost exclusively preferred to model generations, so we exclusively collect inter-variant comparisons.
We randomly sample 5 chapter indices per story in our test set and generate continuations for each model. Participants are shown the story-information and two possible continuations and asked to provide preferences along the above dimensions. We collect one judgement per datapoint (20 datapoints per variant-comparison), and convert the preferences into relative model strengths using a Bradley-Terry model (Bradley & Terry, 1952).
PlotCharacterCreativityDevelopLanguageOverall00111122223333StrengthBBRSFTRL
@@
Figure 2: Bradley Terry relative strength parameters for each variant, by model size. RL-Trained (RL) has highest relative strength in most dimensions, and the effect is more pronounced in 7B models. Table 2 shows win-probabilities for 7B models. Variant shorthands are introduced in Section 7.
@@ 8 Results
SFT model performs poorly on next chapter prediction.   We find significant repetition issues in the chapters generated from the SFT model. For a fair comparison, we automatically truncate clear mode-collapse repetitions (Appendix B), but annotators still strongly prefer all other variants. We hypothesize that SFT performance would be improved by training on a much larger dataset to reduce the effect of overfitting; doing so may also be helpful as a starting point for reasoning models (DeepSeek-AI et al., 2025).
@@
Table 4: We calculate automated measures of length and diversity from the generated chapters across 7B model outputs: # Words (average number of words per chapter), % Unique words within a chapter, % Unseen Trigrams (of all trigrams in a chapter not present in $SI$). Rouge-L (F1 and Precision) is computed against gold standard, reference chapters. We find broadly similar results and uniqueness (Appendix B), indicating that the preferences come from higher quality storytelling instead of longer stories.
@@
We investigate surface-level differences between the generated chapters by comparing their length, word/trigram diversity. While we do find the SFT-based chapters to be significantly shorter and less lexically diverse (Table 4), we find few statistical differences between the other models. This indicates annotators were not biased towards longer/more varied stories, and instead judged based on story-content. Table 9 and Appendix B contain more details.
Perplexity improvement correlates with human judgements.  To evaluate VR-CLI’s Improvement’s utility as a proxy metric for human judgements, we compute the correlation between the pairwise judgement and the Improvement given a reasoning trace and gold next-chapter (Appendix D.1). We find a significant and positive Spearman’s rank correlation of $(\rho=0.33,p=0.01,N=60)$. While this improvement metric cannot be computed at true test time (without a gold continuation), it can be a useful proxy for further evaluating reasoning models on an unlabeled dataset without collecting expensive human judgements.
Reasoning has the most pronounced effect on the scifi genre.  We compare pairwise preferences and average percent-improvement across different genres. We find that reasoning has a particularly strong positive effect on Scifi, and that training further improves the effect. Table 3 shows the percent improvement for Base-Reasoning and RL-Trained variants. See Tables 17 & 18 for pairwise-preferences broken down by model size and genre.
Reasoning improves bigger models more.  To test if these trends are exclusive to 7B models we also perform the same experiments using Qwen 2.5 3B as a base. Relative strength scores are shown in 2(b) and preference probabilities in Table 13; our RL-Trained setting still outperforms the baselines across most dimensions, but the effect is smaller and annotators preferred Baseline-Reasoning in ‘Creativity’ and ‘Language’ dimensions. We find similar genre effects (Scifi still has the highest percent-improvement) and limited lexical differences between settings, more details are in Appendix F. Due to the greater relative strength of RL-Trained 7B variants compared to their non-trained variants, we hypothesize that larger models are better able to learn useful reasoning for this task.
@@ 9 Conclusion
We introduce a novel reward modeling framework, Verifiable Rewards via Completion Likelihood Improvement (VR-CLI), that allows us to optimize a reasoning model to predict plans that increase the likelihood of the next chapter. We train 7B and 3B Qwen models using this framework and GRPO, and show that the resulting story completions are strongly preferred by expert judges over baselines. In the future, we plan to apply our VR-CLI framework to generation tasks whose output cannot be easily verified such as long-form summarization and question answering.
@@ A.1 Book Selection
We selected fiction books based on popularity, critical acclaim, and the availability of chapter summaries on SuperSummary. We excluded all sequels and books based on previous stories (e.g., retellings of myths) to ensure that all relevant information was contained within the book itself. We also excluded collections of short stories, stories featuring existing characters (e.g., a known detective character), and stories relying on visual information (e.g., graphic novels). We also aimed to include a variety of genres: our train/test/validation set all satisfy the following conditions:
@@
    At least one scifi book (that is not also fantasy), one fantasy book (that is not also scifi), and one book that is neither
@@ A.3 Character Sheets
We use the same generation hyper-parameters as the original CHIRON work (Gurung & Lapata, 2024), but adapted to the generation config of Llama 3.3 70B (temperature$=0.6$, top_p$=0.9$).
@@ A.4 Filtering
  • Chapters ($c_{i+1}$) must have $\geq$ 200 words and $\leq$ 5000 words. We also apply the upper limit filter to the previous chapter, to reduce the potential size of the prompt.
@@ A.6 Training
Table 8: Hyper-parameters and their values.
More details for reproducing both setups are available in our code. We base our GRPO-training on (Hu et al., 2024). In both cases we selected the best performing checkpoint via validation performance (percent improvement for GPRO and loss for SFT).
A.7 Lessons from Hyper-parameter Tuning
Due to the high computational cost of our experiments, we were unable to do large hyper-parameter sweeps. We did, however, run many initial tests to gain intuitions on useful hyper-parameters, which we share here to help guide future research in the space.
@@ A.7 Lessons from Hyper-parameter Tuning
Similar to recent results in verifiable reward domains (e.g., math; Lambert et al. 2024), we found that a low KL term (the default parameter in some libraries is 0.05, we use 1e-6) improved performance. We hypothesize that lowering this term encourages more exploration by the policy model, and that our reward definition is robust enough that it doesn’t require significant regularization. However, we do find some cases of significant repetitions in our reasoning traces which may be alleviated by 1) training on a larger dataset and 2) increasing the KL-divergence coefficient.
@@
We found that increasing the number of generations per sample improved performance even when accounting for the additional computational cost. As described earlier the initial average improvement is very low, so we hypothesize that increasing the number of generations increases the likelihood of a non-zero reward, allowing the model to learn from more examples at the beginning of training.
@@ Appendix B Chapter Generation
In order to make model generations more closely comparable to the original chapters and to each other, we bound the length of our next-chapter generations to be between half and 1.5 times the length of the original (in tokens): $[0.5\cdot|c_{i+1}|,1.5\cdot|c_{i+1}|]$
@@
For fairness we do not apply these rules to reasoning traces (to simulate a hands-off case using reasoning), although they do occasionally also suffer from mode-collapse and repetitions.
@@
Table 4 contains automated measures of length and diversity, across our chapter generation variants. The biggest differences come from the SFT-based chapters, which produce shorter and less diverse chapters. In contrast, the differences between our other variants are fairly small - performing a one-way ANOVA test between each of our variants (one test per metric e.g., # Words) does not produce any statistically significant results $(p>0.05)$, with the exception of Rouge-L Precision. As our automated metrics largely imply lexically similar stories across variants we believe that our annotators were not swayed in their judgments by differences in length or lexical diversity.
@@
We further break down Rouge-L Precision in Table 10 by using different pieces of story-information as reference. We find that both generated chapters and reasoning traces draw heavily from all pieces of story information, but reasoning traces seem to more heavily use the Next Chapter Synopsis and Character Sheets. Trained reasoning traces also tend to have higher Rouge-L Precision than un-trained reasoning, across 3B and 7B variants and story-information categories.
@@
Table 9: We calculate automated measures of length and diversity from the generated chapters across our variants: # Words (average number of words per chapter), % Unique Words (percentage of unique words within a chapter), % Unseen Trigrams (percent of all trigrams in a chapter that were not present in $SI$). We find broadly similar lengths and uniqueness, indicating that the preferences come from higher quality storytelling instead of longer stories. We also report metrics for the reasoning traces ‘REAS’, which show greater variability.
@@
Table 10: Rouge-L Precision for the generated next-chapters and the reasoning traces, using different pieces of the story-information $SI$ as reference. We find that the reasoning traces contain more overlap with all pieces of the story-information, but that there doesn’t seem to be an obvious correlation between specific pieces of the story-information and model variant success. ‘Prev. Chap.’ refers to the previous chapter, ‘Prior Sum.’ to the summary of the previous story, ‘Plot Sketch’ to the global plot sketch, ‘CSheets’ to the character sheets, and ‘Next Chap. Syn.’ to the synopsis of the next-chapter.

### v2

@@ 6 Reinforcement Learning for Next-Chapter Prediction
[…]
Applying VR-CLI to our task implies that plans that induce a higher likelihood of the true next chapter will also produce better generated chapters. We validate this assumption in Section 8, but also believe it aligns with our intuition concerning useful plans. For example, plans that correctly predict plot events, character details, or writing style in the true next chapter should increase its likelihood, and should also be a useful basis for generating a novel chapter. We provide example aligned excerpts between plans and next-chapters in Table 19.
We define our thresholded reward as:
[…]
@@
Our task and reward setup are relatively RL-algorithm agnostic; one could use an offline reasoning-reward dataset or any online policy learning algorithm that works with a verifiable reward. We hypothesize that online methods are preferable for NCP as reasoning styles change over training, and preliminary tests show low reward prior to training: initially, average improvement is -0.06% with Qwen 2.5 7B.
@@ 7.1 Model Comparisons
[…]
@@
[…]
@@
We fine-tune our story-generator on the NCP task with the SFT objective on next chapters. The generator predicts chapters like Base: $\hat{c}_{i}\leftarrow\pi^{\mathcal{G}}_{\theta}(SI_{i})$.
@@ 7.2 Human Evaluation
We evaluate the generated story continuations by eliciting pairwise preferences following evidence suggesting that relative judgments can be more reliable than absolute ones (Louviere et al., 2015; Stewart et al., 2005; Liu et al., 2024). We draw on the criteria outlined in previous work (Huot et al., 2025; Chakrabarty et al., 2024; Chhun et al., 2022) to evaluate the continuations along the following dimensions: (1) Plot: Does it exhibit events and turns that move the plot forward logically? (2) Creativity: Does the continuation have engaging characters, themes, and imagery, and avoid overly cliched characters and storylines? (3) Development:   Does it introduce characters and settings with appropriate levels of detail and complexity? (4) Language Use:   Is the language varied and rich, exhibiting rhetorical, linguistic, and literary devices? (5) Characters: Does it feature believable and conceptually consistent characters, including reasonable character arcs and development? (6) Overall Preference: Which of the two continuations did you prefer? Full instructions and recruitment details are in Appendix C. Annotators were recruited through Prolific if their primary language was English and their employment role was in creative writing.
To validate our annotation procedure, we compute Fleiss’ kappa across a dataset of Base vs. Base-Reasoning comparisons. We find fair agreement across annotation dimensions, with the highest agreement in Creativity and Overall Quality. More details are in Appendix D. Initial experiments showed gold-continuations were almost exclusively preferred to model generations, so we exclusively collect inter-variant comparisons.
We randomly sample 5 chapter indices per story in our test set and generate continuations for each model. Participants are shown the story information and two possible continuations and asked to provide preferences along the above dimensions. We collect one judgment per datapoint (20 datapoints per variant-comparison), and convert the preferences into relative model strengths using a Bradley-Terry model (Bradley & Terry, 1952).
PlotCharacterCreativityDevelopLanguageOverall0112233StrengthBBRSFTRL
@@
Figure 2: Bradley Terry relative strength parameters for each variant, by model size. RL-Trained (RL) has the highest relative strength in most dimensions, and the effect is more pronounced in 7B models. Table 2 shows win-probabilities for 7B models. Variant shorthands are introduced in Section 7.
@@ 8 Results
SFT model performs poorly on next chapter prediction.   We find significant repetition issues in the chapters generated from the SFT model. For a fair comparison, we automatically truncate clear mode-collapse repetitions (Appendix B), but annotators still strongly prefer all other variants. We hypothesize that SFT performance would be improved by training on a much larger dataset to reduce the effect of overfitting; doing so may also be helpful as a starting point for reasoning models (DeepSeek-AI et al., 2025). We therefore believe that our method is more effective in low-resource settings, but can be used in concert with SFT approaches when more data is available.
@@
Table 4: We calculate automated measures of length and diversity from the generated chapters across 7B model outputs: # Words (average number of words per chapter), % Unique words within a chapter, % Unseen Trigrams (% of all trigrams in a chapter not present in SISI). Rouge-L (F1 and Precision) is computed against the gold standard, reference chapters. We find broadly similar results and uniqueness (Appendix B), indicating that the preferences come from higher quality storytelling instead of longer stories.
@@
We investigate surface-level differences between the generated chapters by comparing their length, word/trigram diversity. While we do find the SFT-based chapters to be significantly shorter and less lexically diverse (Table 4), we find few statistical differences between the other models. This indicates that annotators were not biased towards longer/more varied stories and instead judged based on story content. Table 9 and Appendix B contain more details.
Perplexity improvement correlates with human judgments.  To evaluate VR-CLI’s Improvement’s utility as a proxy metric for human judgments, we compute the correlation between the pairwise judgment and the Improvement given a reasoning trace and gold next-chapter (Appendix D.1). We find a significant and positive Spearman’s rank correlation of $(\rho=0.33,p=0.01,N=60)$. While this improvement metric cannot be computed at true test time (without a gold continuation), it can be a useful proxy for further evaluating reasoning models on an unlabeled dataset without collecting expensive human judgments.
Reasoning has the most pronounced effect on the Sci-Fi genre.  We compare pairwise preferences and average percent improvement across different genres. We find that reasoning has a particularly strong positive effect on Sci-Fi, and that training further improves the effect. Table 3 shows the percent improvement for Base-Reasoning and RL-Trained variants. See Tables 17 & 18 for pairwise preferences broken down by model size and genre.
Reasoning improves bigger models more.  To test if these trends are exclusive to 7B models, we also perform the same experiments using Qwen 2.5 3B as a base. Relative strength scores are shown in 2(b) and preference probabilities in Table 13; our RL-Trained setting still outperforms the baselines across most dimensions, but the effect is smaller and annotators preferred Baseline-Reasoning in ‘Creativity’ and ‘Language’ dimensions. We find similar genre effects (Sci-fi still has the highest percent-improvement) and limited lexical differences between settings; more details are in Appendix F. Due to the greater relative strength of RL-Trained 7B variants compared to their non-trained variants, we hypothesize that larger models are better able to learn useful reasoning for this task.
@@ 9 Conclusion
We introduce a novel reward modeling framework, Verifiable Rewards via Completion Likelihood Improvement (VR-CLI), that allows us to optimize a reasoning model to predict plans that increase the likelihood of the next chapter. We train 7B and 3B Qwen models using this framework and GRPO, and show that the resulting story completions are strongly preferred by expert judges over baselines. In the future, we plan to apply our VR-CLI framework to generation tasks whose output cannot be easily verified, such as long-form summarization and question answering.
Acknowledgments
We thank the anonymous reviewers for their constructive feedback. We gratefully acknowledge the support of the UK Engineering and Physical Sciences Research Council (grant EP/W002876/1). We would also like to thank Kolya Malkin for their advice and support on this work.
@@ A.1 Book Selection
We selected fiction books based on popularity, critical acclaim, and the availability of chapter summaries on SuperSummary. We excluded all sequels and books based on previous stories (e.g., retellings of myths) to ensure that all relevant information was contained within the book itself. We also excluded collections of short stories, stories featuring existing characters (e.g., a known detective character), and stories relying on visual information (e.g., graphic novels). We also aimed to include a variety of genres; our train/test/validation set all satisfy the following conditions:
@@
    At least one Sci-Fi book (that is not also fantasy), one fantasy book (that is not also Sci-Fi), and one book that is neither
@@ A.3 Character Sheets
We use the same generation hyper-parameters as the original CHIRON work (Gurung & Lapata, 2024), but adapted to the generation config of Llama 3.3 70B (temperature=0.6=0.6, top_p=0.9=0.9).
@@ A.4 Filtering
  • Chapters ($c_{i+1}$) must have $\geq$ 200 words and $\leq$ 5000 words. We also apply the upper limit filter to the previous chapter to reduce the potential size of the prompt.
@@ A.6 Training
Epochs 10
Table 8: Hyper-parameters and their values. After training, models were selected by the best-performing checkpoint on validation loss/reward.
More details for reproducing both setups are available in our code. We base our GRPO-training on (Hu et al., 2024b). In both cases we selected the best performing checkpoint via validation performance (percent improvement for GPRO and loss for SFT).
A.7 SFT Sweeps
We ran small sweeps to test different SFT hyper-parameters, selected from previous work comparing SFT to RL methods. We evaluated learning rates of {2e-5, 5e-7, 5e-6, 4e-5}, and amongst largely similar results found 2e-5 to perform the best by validation loss. Future work could likely optimize further, but we believe that the largest barrier to high-quality SFT performance is dataset size.
A.8 Reward Formulation Ablations (VR-CLI Variants)
Recall that for this work, we defined our perplexity-based Improvement measure as:
$I_{\pi^{\mathcal{G}}}(x,y,a)=[\frac{PPL_{\pi^{\mathcal{G}}}(y|x)-PPL_{\pi^{\mathcal{G}}}(y|x,a)}{PPL_{\pi^{\mathcal{G}}}(y|x)}]\times 100=[1-\frac{PPL_{\pi^{\mathcal{G}}}(y|x,a)}{PPL_{\pi^{\mathcal{G}}}(y|x)}]\times 100$
and our reward as a piecewise function (although we propose other reward formulations):
[…]
We also ran ablations to test different formulations of our reward, within the same VR-CLI paradigm. Instead of costly human annotations, we evaluate these different methods via their average percent (perplexity) improvement on our test set, using reasoning produced after RL-training. We sample 5 plans to reduce variance. Our “RL-Trained” method, with the previously described improvement and reward formulation, produces an average percent-improvement of $0.884$.
NLL for PPL   Our first ablation keeps everything the same, but switches out perplexity for log-likelihood in the Improvement calculation. We find a new percent-improvement of $0.485\%$, about $45\%$ worse than our method. We noticed a lack of convergence with the same hyper-parameters, so we re-ran training for 30 epochs instead of 20, resulting in a percent improvement of $0.627\%$, still $30\%$ worse.
Unbounded NLL Improvement  Our second ablation tests a much simpler version of our reward calculation: the Improvement calculation is reduced to the log-likelihood (without a baseline calculation, and the ‘unbounded’ reward is set equal to the improvement.
$I_{\pi^{\mathcal{G}}}(x,y,a)=\log P_{\pi^{\mathcal{G}}}(y|x,a)$
[…]
We find slightly better performance with this simpler reward compared to our default setting, with an average percent-improvement of $1.031\%$ (about $16\%$ higher than our method). This differs from our initial experiments with 3B models (which may have worse reward distributions), which led us to develop the piecewise reward formulation. However, as described in Section 5, one downside to this approach is the worse interpretability during training, as 1) it can be unclear whether performance is better without reasoning, and 2) some samples may dominate the logged metrics.
Unbounded PPL Improvement  Interestingly, when we performed the same experiment with perplexity ($I_{\pi^{\mathcal{G}}}(x,y,a)=-PPL_{\pi^{\mathcal{G}}}(y|x,a)$) instead of log-likelihood our performance dropped to an average of $0.760$.
An interesting avenue for future work is a detailed empirical comparison of these VR-CLI formulations on creative writing tasks and beyond.
A.9 Lessons from Hyper-parameter Tuning
Due to the high computational cost of our experiments, we were unable to do large hyper-parameter sweeps. We did, however, run many initial tests to gain insight on useful hyper-parameters, which we share here to help guide future research in the space.
@@ A.9 Lessons from Hyper-parameter Tuning
Similar to recent results in verifiable reward domains (e.g., math; Lambert et al. 2024), we found that a low KL term (the default parameter in some libraries is 0.05; we use 1e-6) improved performance. We hypothesize that lowering this term encourages more exploration by the policy model, and that our reward definition is robust enough that it doesn’t require significant regularization. However, we do find some cases of significant repetitions in our reasoning traces, which may be alleviated by 1) training on a larger dataset and 2) increasing the KL-divergence coefficient.
@@
We found that increasing the number of generations per sample improved performance even when accounting for the additional computational cost. As described earlier, the initial average improvement is very low, so we hypothesize that increasing the number of generations increases the likelihood of a non-zero reward, allowing the model to learn from more examples at the beginning of training.
@@ Appendix B Chapter Generation
To make model generations more closely comparable to the original chapters and each other, we bound the length of our next-chapter generations to be between half and 1.5 times the length of the original (in tokens): $0.5\cdot|c_{i+1}|,1.5\cdot|c_{i+1}|]$
@@
For fairness, we do not apply these rules to reasoning traces (to simulate a hands-off case using reasoning), although they do occasionally also suffer from mode-collapse and repetitions.
@@
Table 4 contains automated measures of length and diversity, across our chapter generation variants. The biggest differences come from the SFT-based chapters, which produce shorter and less diverse chapters. In contrast, the differences between our other variants are fairly small - performing a one-way ANOVA test between each of our variants (one test per metric, e.g., # Words) does not produce any statistically significant results $(p>0.05)$, with the exception of Rouge-L Precision. As our automated metrics largely imply lexically similar stories across variants, we believe that our annotators were not swayed in their judgments by differences in length or lexical diversity.
@@
We further break down Rouge-L Precision in Table 10 by using different pieces of story information as reference. We find that both generated chapters and reasoning traces draw heavily from all pieces of story information, but reasoning traces seem to more heavily use the Next Chapter Synopsis and Character Sheets. Trained reasoning traces also tend to have higher Rouge-L Precision than untrained reasoning, across 3B and 7B variants and story-information categories.
@@
Table 9: We calculate automated measures of length and diversity from the generated chapters across our variants: # Words (average number of words per chapter), % Unique Words (percentage of unique words within a chapter), % Unseen Trigrams (percent of all trigrams in a chapter that were not present in SISI). We find broadly similar lengths and uniqueness, indicating that the preferences come from higher-quality storytelling instead of longer stories. We also report metrics for the reasoning traces ‘REAS’, which show greater variability.
@@
Table 10: Rouge-L Precision for the generated next-chapters and the reasoning traces, using different pieces of the story-information SISI as reference. We find that the reasoning traces contain more overlap with all pieces of the story information, but that there doesn’t seem to be an obvious correlation between specific pieces of the story information and model variant success. ‘Prev. Chap.’ refers to the previous chapter, ‘Prior Sum.’ to the summary of the previous story, ‘Plot Sketch’ to the global plot sketch, ‘CSheets’ to the character sheets, and ‘Next Chap. Syn.’ to the synopsis of the next-chapter.